# Descent-Guided Policy Gradient for Scalable Cooperative Multi-Agent Learning

### v1

@@ 3.3 Implementation
The policy and critic are then updated using standard PPO objectives with the modified advantages. The computational overhead is negligible: the reference trajectory $\{\tilde{\boldsymbol{x}}_{t}\}$ is computed once per batch from the analytical model, and in the common case where actions affect the system state linearly, the guidance coefficient reduces to a simple inner product computable in $\mathcal{O}(d)$ time.
@@ 4 Theoretical Guarantees
In this section, we establish the theoretical foundations of DG-PG. Our central result is that the DG-PG gradient estimator achieves agent-independent variance (Theorem 4.2), breaking the $\Theta(N)$ scaling barrier of standard policy gradients. We complement this with two supporting guarantees: (1) consistency (Theorem 4.1), ensuring that the augmented objective preserves the stationary points of the original game; and (2) sample complexity (Theorem 4.3), showing that the variance reduction yields agent-independent convergence rates under standard regularity assumptions (the complete dependency structure is given in Table 3, Appendix B).
4.1 Consistency (Nash Invariance)
A primary concern when optimizing the augmented objective $J_{\alpha}=(1-\alpha)J-\alpha\mathcal{G}$ is whether it biases the optimization landscape. We prove that under our assumptions, the stationary points of the original objective are preserved.
@@ 4.1 Consistency (Nash Invariance)
Let $\boldsymbol{\theta}^{*}$ be a stationary point of the original cooperative objective, satisfying $\nabla_{\boldsymbol{\theta}}J(\boldsymbol{\theta}^{*})=\mathbf{0}$. Under Assumptions 3.13.2, $\boldsymbol{\theta}^{*}$ remains a stationary point of the augmented objective:
$\nabla_{\boldsymbol{\theta}}J(\boldsymbol{\theta}^{*})=\mathbf{0}\implies\nabla_{\boldsymbol{\theta}}J_{\alpha}(\boldsymbol{\theta}^{*})=\mathbf{0},\quad\forall\alpha\in(0,1).$ (12)
This result guarantees the safety of DG-PG: optimizing $J_{\alpha}$ does not introduce spurious equilibria, ensuring that any optimal policy for the original objective remains optimal under the augmented objective.
@@
If $\nabla_{\boldsymbol{\theta}}\mathcal{G}(\boldsymbol{\theta}^{*})\neq\mathbf{0}$ at a stationary point of $J$, then $\boldsymbol{x}(\boldsymbol{\theta}^{*})\neq\tilde{\boldsymbol{x}}$ (otherwise $\mathcal{G}$ is minimized). Assumption 3.2 then forces $\nabla_{\boldsymbol{x}}J\neq\mathbf{0}$, and sufficient policy expressiveness lifts this to $\nabla_{\boldsymbol{\theta}}J\neq\mathbf{0}$, contradicting stationarity. The detailed proof is in Appendix E. ∎
@@ 4.2 Variance Reduction
Having established that the augmented objective preserves the stationary points of $J$ (Theorem 4.1), we now formalize the variance properties of the DG-PG gradient estimator. The key structural difference between the two gradient terms lies in their coefficients: $G_{t}$ depends on all $N$ agents’ stochastic actions, whereas the guidance coefficient $\langle\boldsymbol{x}_{t}-\tilde{\boldsymbol{x}}_{t},\boldsymbol{z}_{t}^{i}\rangle$ depends only on observed and agent-local quantities (Section 3.2). This yields fundamentally different variance scaling: $\mathrm{Var}(\hat{g}_{J}^{i})=\Theta(N)$, while $\mathrm{Var}(\hat{g}_{\mathcal{G}}^{i})=\mathcal{O}(1)$, independent of $N$.
Since the DG-PG estimator is a convex combination of $\hat{g}_{J}^{i}$ and $\hat{g}_{\mathcal{G}}^{i}$, the combined variance depends on their correlation $\rho=\mathrm{Corr}(\hat{g}_{J}^{i},\hat{g}_{\mathcal{G}}^{i})$. Under Assumption 3.2, the two gradients are positively correlated ($\rho>0$) since both directions improve $J$; they are imperfectly correlated ($\rho<1$) because $\hat{g}_{J}^{i}$ additionally carries cross-agent noise absent from $\hat{g}_{\mathcal{G}}^{i}$. The following theorem characterizes the optimal variance as a function of $\rho$ and the individual variances.
@@
Under Assumptions 3.13.2, let $\sigma_{J}^{2}\triangleq\mathrm{Var}(\hat{g}_{J}^{i})$, $\sigma_{\mathcal{G}}^{2}\triangleq\mathrm{Var}(\hat{g}_{\mathcal{G}}^{i})$, and $\rho\triangleq\mathrm{Corr}(\hat{g}_{J}^{i},\hat{g}_{\mathcal{G}}^{i})\in(0,1)$. The optimally weighted combined estimator achieves:
@@
Since $\sigma_{J}^{2}=\Theta(N)$ and $\sigma_{\mathcal{G}}^{2}=\mathcal{O}(1)$ (established above), the minimum variance converges to $\sigma_{\mathcal{G}}^{2}(1-\rho^{2})=\mathcal{O}(1)$ as $N\to\infty$, independent of the number of agents.
@@
The variance of $\hat{g}_{\mathrm{DG}}^{i}=(1-\alpha)\hat{g}_{J}^{i}-\alpha\hat{g}_{\mathcal{G}}^{i}$ is a convex quadratic in $\alpha$. Minimizing in closed form and substituting $\sigma_{J}^{2}=\Theta(N)$, $\sigma_{\mathcal{G}}^{2}=\mathcal{O}(1)$ yields the $\mathcal{O}(1)$ bound. Since $\sigma_{\mathcal{G}}^{2}/\sigma_{J}^{2}=\mathcal{O}(1/N)$, the optimal mixing coefficient satisfies $\alpha^{*}\to 1$ as $N\to\infty$, indicating that large systems should rely primarily on the guidance signal. The full derivation is provided in Appendix F. ∎
The role of $\alpha$. Although $\alpha^{*}\to 1$ is optimal in the large-$N$ limit, two practical factors prevent setting $\alpha$ close to $1$. First, the guidance variance $\sigma_{\mathcal{G}}^{2}$ persists—the score function $\nabla\log\pi^{i}$ introduces irreducible stochasticity that does not vanish with NN. Second, in many practical domains, Assumptions 3.13.2 may not hold exactly, in which case the augmented objective $J_{\alpha}$ may no longer share the same optima as $J$. Consequently, the choice of $\alpha$ must balance variance reduction against fidelity to the original objective. We investigate the sensitivity to $\alpha$ and the effect of scheduling strategies empirically in Section 6.
@@ 4.3 Convergence Rate
Assume $J_{\alpha}$ is $L$-smooth and satisfies the Polyak-Łojasiewicz (PL) condition with parameter $\mu>0$ (see Assumption G.1 in Appendix G). Under the variance bound from Theorem 4.2, DG-PG achieves an $\epsilon$-suboptimal solution in:
$T=\mathcal{O}\left(\frac{L}{\mu\epsilon}\right)$ (14)
iterations. This bound is independent of the number of agents $N$.
In contrast, standard multi-agent policy gradient methods require $T=\mathcal{O}(N/\epsilon)$ iterations to compensate for $\Theta(N)$ variance—the required number of iterations grows linearly in $N$, whereas DG-PG’s iteration count remains unchanged.
@@
Under Assumption G.1, standard stochastic gradient ascent converges at a rate governed by the gradient variance $\sigma^{2}$. Substituting $\sigma^{2}=\mathcal{O}(1)$ from Theorem 4.2 yields the NN-independent bound. The full derivation is provided in Appendix G. ∎
Summary. Theorems 4.14.3 collectively establish that DG-PG is (1) scalable—variance does not grow with NN; (2) safe—it preserves optimal solutions; and (3) efficient—sample complexity is agent-independent.
@@ 5 Related Work
Credit Assignment in Cooperative MARL. Credit decomposition methods attempt to extract per-agent contributions from the shared return. Difference Rewards (Wolpert and Tumer, 2001) isolate each agent’s marginal contribution via counterfactual simulation, but require access to the reward function or a simulator. Value decomposition methods—VDN (Sunehag et al., 2017), QMIX (Rashid et al., 2020), QPLEX (Wang et al., 2021a)—factorize the joint Q-function into per-agent terms under structural assumptions (additivity, monotonicity), avoiding gradient estimation but restricting the class of representable objectives. COMA (Foerster et al., 2018) trains a centralized critic to compute counterfactual baselines that marginalize each agent’s action, but the critic must condition on the joint action space and becomes difficult to train as NN grows. All these methods ultimately derive their training signal from the shared return, so the cross-agent noise that grows with NN persists.
@@
Variance Reduction in Multi-Agent Policy Gradients. Standard variance reduction techniques—value function baselines (Sutton and Barto, 1998) and GAE (Schulman et al., 2016)—primarily reduce temporal variance from an agent’s own stochasticity. In the multi-agent setting, HAPPO (Kuba et al., 2022) reduces cross-agent noise by updating agents sequentially, conditioning each agent’s update on the previous agents’ new policies. This is effective but sacrifices parallel execution, as each agent must wait for all preceding updates. Agent-specific baselines can reduce the self-exploration component of variance but do not explicitly target the cross-agent noise that dominates in large-scale settings. Mean-Field MARL (Yang et al., 2018) takes a different approach, replacing the $N$-agent interaction with a population-level approximation that simplifies each agent’s optimization problem. This achieves scalability but requires agent homogeneity, precluding application to heterogeneous settings.
Reward Shaping. Like DG-PG, Potential-Based Reward Shaping (PBRS) (Ng et al., 1999) augments the objective while preserving optimal policies, so a natural question is how the two differ. PBRS reshapes the reward via a potential function to provide denser temporal feedback, accelerating single-agent exploration; it does not decompose across agents and therefore does not target cross-agent noise. DG-PG’s augmentation serves a different purpose: its gradient decomposes into per-agent terms that can each be computed deterministically from the system state and each agent’s own action, directly reducing the multi-agent noise in gradient estimation.
@@
Structural Priors in RL. Guided Policy Search (Levine and Koltun, 2013) and Residual RL (Johannink et al., 2019) incorporate domain knowledge by using pre-computed controllers as base policies or supervisory signals, but this couples the learned policy to the quality of the prior and can bias the solution. In MARL, graph-based communication (Jiang et al., 2020) and role decomposition (Wang et al., 2021b) use structural knowledge as architectural inductive biases to reduce the search space, though they do not directly address the cross-agent noise in gradient estimation. DG-PG uses analytical models differently: to construct noise-free per-agent guidance gradients that reduce gradient estimate variance without altering the optimization landscape. A systematic comparison is provided in Table 2 (Appendix A).
@@ 6.1 Setup
Environment. We evaluate on a heterogeneous cloud scheduling simulator with NN servers whose configurations are based on AWS instance types. The environment combines three sources of complexity—heterogeneous server efficiency, bimodal workloads with heavy tails, and non-stationary arrivals—that jointly violate the stationarity assumptions of the queueing-theoretic prior, testing robustness to model mismatch. The reward balances two competing objectives: latency (queue minimization) and energy efficiency (preferring high-$\eta_{\text{CPU}}$ servers), which cannot be simultaneously optimized by simple heuristics. Figure 1 illustrates these characteristics. We evaluate at scales $N\in\{5,10,20,50,100,200\}$; for each scale, we pre-generate 200 training scenarios and 40 held-out test scenarios with randomized server configurations and workload patterns. Full specifications are in Appendix H.
@@ 6.2 Hyperparameter Selection
The guidance weight $\alpha\in(0,1)$ controls the bias-variance trade-off in DG-PG: larger $\alpha$ reduces gradient variance by relying more heavily on the analytical prior, while smaller $\alpha$ maintains closer alignment with the true reward signal. We conduct a systematic grid search over $\alpha\in\{0.2,0.4,0.6,0.8\}$ at $N=50$ using NN models, with validation at $N=20$.
@@ Results.
Figure 2 shows training curves at N=50N{=}50 for all tested $\alpha$ values. We observe a clear trade-off: larger $\alpha$ accelerates early convergence, while smaller $\alpha$ achieves marginally superior final performance. By episode 200, $\alpha{=}0.2$ achieves the best mean training reward of 40.3-40.3, while $\alpha{=}0.8$ reaches 40.8-40.8—a gap of less than 1%. This trend is consistent at N=20N{=}20, where $\alpha{=}0.2$ achieves 39.9-39.9. The narrow spread across $\alpha$ values demonstrates that DG-PG is robust to the choice of $\alpha$: the RL gradient compensates for model inaccuracy while benefiting from variance reduction at any tested setting.

### v2

@@ 3.3 Implementation
The policy and critic are then updated using standard PPO objectives with the modified advantages. The computational overhead is minimal, since the reference trajectory $\{\tilde{\boldsymbol{x}}_{t}\}$ is computed once per batch from the analytical model, and each guidance coefficient amounts to an $\mathcal{O}(m)$ inner product per agent.
@@ 4 Theoretical Guarantees
In this section, we establish the theoretical foundations of DG-PG. Our central result is that the DG-PG gradient estimator achieves agent-independent variance (Theorem 4.2), breaking the $\Theta(N)$ scaling barrier of standard policy gradients. We complement this with two supporting guarantees: (1) consistency (Theorem 4.1), ensuring that DG-PG preserves the Nash equilibria of the original game; and (2) sample complexity (Theorem 4.3), showing that the variance reduction yields agent-independent convergence rates under standard regularity assumptions (the complete dependency structure is given in Table 3, Appendix B).
4.1 Nash Invariance
A primary concern when augmenting the policy gradient is whether the additional guidance term shifts the Nash equilibria of the original cooperative game. We prove that under our assumptions, $\nabla_{\theta^{i}}J_{\alpha}$ vanishes at every Nash equilibrium of $J$.
@@ 4.1 Nash Invariance
Let $\boldsymbol{\theta}^{*}$ be a Nash equilibrium of the original cooperative game, i.e., $\nabla_{\theta^{i}}J(\boldsymbol{\theta}^{*})=\mathbf{0}$ for all $i\in\mathcal{N}$. Under Assumptions 3.13.2, the DG-PG gradient also vanishes at $\boldsymbol{\theta}^{*}$:
$\nabla_{\theta^{i}}J_{\alpha}(\boldsymbol{\theta}^{*})=\mathbf{0},\quad\forall\alpha\in(0,1),\;\forall i\in\mathcal{N}.$ (15)
This result guarantees the safety of DG-PG, as the augmentation of the guidance term does not alter the Nash equilibria of the original game.
@@
We first show by contradiction that $\nabla_{\theta^{i}}\mathcal{G}(\boldsymbol{\theta}^{*})=\mathbf{0}$ at any Nash equilibrium $\boldsymbol{\theta}^{*}$ (Lemma E.1). Since both $\nabla_{\theta^{i}}J$ and $\nabla_{\theta^{i}}\mathcal{G}$ vanish at $\boldsymbol{\theta}^{*}$, the DG-PG gradient $\nabla_{\theta^{i}}J_{\alpha}=(1-\alpha)\cdot\mathbf{0}-\alpha\cdot\mathbf{0}=\mathbf{0}$. The detailed proof is in Appendix E. ∎
@@ 4.2 Variance Reduction
Having established that DG-PG preserves the Nash equilibria of $J$ (Theorem 4.1), we now formalize the variance properties of the DG-PG gradient estimator. The return $R_{t}$ in the standard policy gradient aggregates all $N$ agents’ stochastic actions, so the per-agent gradient variance $\mathrm{Var}(\hat{g}_{J}^{i})=\Theta(N)$. In contrast, the guidance coefficient $\langle\boldsymbol{x}_{t}-\tilde{\boldsymbol{x}}_{t},\boldsymbol{z}_{t}^{i}\rangle$ depends only on agent-local quantities (Section 3.2), yielding guidance gradient variance $\mathrm{Var}(\hat{g}_{\mathcal{G}}^{i})=\mathcal{O}(1)$ independent of $N$.
Since the DG-PG estimator combines $\hat{g}_{J}^{i}$ and $\hat{g}_{\mathcal{G}}^{i}$, the combined variance depends on their correlation $\rho=\mathrm{Corr}(\hat{g}_{J}^{i},\hat{g}_{\mathcal{G}}^{i})$. Under Assumption 3.2, improving $J$ reduces the deviation $\mathcal{G}$, so $\hat{g}_{J}^{i}$ and $\hat{g}_{\mathcal{G}}^{i}$ are negatively correlated ($\rho<0$). The following theorem characterizes the optimal variance as a function of $\rho$ and the variances of two gradient estimators.
@@
Under Assumptions 3.13.2, let $\sigma_{J}^{2}\triangleq\mathrm{Var}(\hat{g}_{J}^{i})$, $\sigma_{\mathcal{G}}^{2}\triangleq\mathrm{Var}(\hat{g}_{\mathcal{G}}^{i})$, and $\rho\triangleq\mathrm{Corr}(\hat{g}_{J}^{i},\hat{g}_{\mathcal{G}}^{i})\in(-1,0)$. The optimally weighted combined estimator achieves:
@@
Since $\sigma_{J}^{2}=\Theta(N)$ and $\sigma_{\mathcal{G}}^{2}=\mathcal{O}(1)$ (established above), the minimum variance satisfies $\min_{\alpha}\mathrm{Var}(\hat{g}_{\mathrm{DG}}^{i})=\mathcal{O}(1)$ for all $N$, independent of the number of agents.
@@
The combined variance $\sigma_{\mathrm{DG}}^{2}(\alpha)$ is a convex quadratic in $\alpha$ with cross-term $\rho\,\sigma_{J}\sigma_{\mathcal{G}}$. Minimizing in closed form yields
$\sigma_{\mathrm{DG},\min}^{2}=\sigma_{\mathcal{G}}^{2}(1-\rho^{2})\,/\,({\sigma_{J}^{2}+\sigma_{\mathcal{G}}^{2}+2\rho\,\sigma_{J}\sigma_{\mathcal{G}}}).$
Substituting $\sigma_{J}^{2}=\Theta(N)$ and $\sigma_{\mathcal{G}}^{2}=\mathcal{O}(1)$, the denominator is dominated by $\sigma_{J}^{2}$ and is thus bounded away from zero, while the numerator is $\mathcal{O}(1)$, giving $\sigma_{\mathrm{DG},\min}^{2}=\mathcal{O}(1)$ for all $N$. The corresponding optimal weight $\alpha^{*}$ satisfies $\alpha^{*}\to 1$ as NN grows, indicating that larger systems rely more on the guidance signal. The full derivation is provided in Appendix F. ∎
The role of $\alpha$. Although $\alpha^{*}\to 1$ is optimal in the large-$N$ limit, two practical factors prevent setting $\alpha$ close to $1$. First, since the score function $\nabla\log\pi^{i}$ introduces irreducible stochasticity that does not vanish with $N$, the guidance variance $\sigma_{\mathcal{G}}^{2}$ persists. Second, in many practical domains, Assumptions 3.13.2 may not hold exactly, in which case the Nash invariance guarantee (Theorem 4.1) may no longer hold. Consequently, the choice of $\alpha$ must balance variance reduction against fidelity to the original game. We investigate the sensitivity to $\alpha$ and the effect of scheduling strategies empirically in Section 6.
@@ 4.3 Convergence Rate
Assume $J$ is $L$-smooth and satisfies the Polyak-Łojasiewicz (PL) condition with PL constant $\mu\in(0,L]$ (see Assumption G.1 in Appendix G). Under the variance bound from Theorem 4.2, DG-PG achieves an $\epsilon$-optimal policy in:
$T=\mathcal{O}\left(\frac{L}{\mu^{2}\epsilon}\log\frac{1}{\epsilon}\right)$ (17)
iterations. This bound is independent of the number of agents NN.
This result establishes that DG-PG’s convergence rate is independent of the number of agents, so adding more agents does not require additional training iterations. In contrast, standard multi-agent policy gradient methods require $T=\mathcal{O}(N/\epsilon)$ iterations to compensate for $\Theta(N)$ variance.
@@
$L$-smoothness bounds the per-iteration descent in terms of $\|\nabla J\|^{2}$ and the gradient variance $\sigma_{\mathrm{DG}}^{2}$. The PL condition then converts $\|\nabla J\|^{2}$ into the optimality gap $J^{*}-J(\boldsymbol{\theta}_{k})$. Substituting $\sigma_{\mathrm{DG}}^{2}=\mathcal{O}(1)$ from Theorem 4.2 yields the $N$-independent bound. The full derivation is provided in Appendix G. ∎
Summary. Theorems 4.14.3 collectively establish that DG-PG is (1) scalable: variance does not grow with NN; (2) safe: optimal solutions are preserved; and (3) efficient: sample complexity is agent-independent.
@@ 5 Related Work
Credit Assignment in Cooperative MARL. Credit decomposition methods attempt to extract per-agent contributions from the shared return. Difference Rewards (Wolpert and Tumer, 2001) isolate each agent’s marginal contribution via counterfactual simulation, but require access to the reward function or a simulator. Value decomposition methods, such as VDN (Sunehag et al., 2017), QMIX (Rashid et al., 2020), and QPLEX (Wang et al., 2021), factorize the joint Q-function into per-agent terms under structural assumptions (additivity, monotonicity), avoiding gradient estimation but restricting the class of representable objectives. COMA (Foerster et al., 2018) trains a centralized critic to compute counterfactual baselines that marginalize each agent’s action, but the critic must condition on the joint action space and becomes difficult to train as NN grows. All these methods still derive their training signal from the shared return, so the cross-agent noise that grows with NN persists.
@@
Variance Reduction in Multi-Agent Policy Gradients. Standard variance reduction techniques, including value function baselines (Sutton and Barto, 1998) and GAE (Schulman et al., 2016), primarily reduce temporal variance from an agent’s own stochasticity. In the multi-agent setting, HAPPO (Kuba et al., 2022) reduces cross-agent noise by updating agents sequentially, conditioning each agent’s update on the previous agents’ new policies. This is effective but sacrifices parallel execution, as each agent must wait for all preceding updates. Agent-specific baselines can reduce the self-exploration component of variance but do not explicitly target the cross-agent noise that dominates in large-scale settings. Mean-Field MARL (Yang et al., 2018) takes a different approach, replacing the $N$-agent interaction with a population-level approximation that simplifies each agent’s optimization problem. This achieves scalability but requires agent homogeneity, precluding application to heterogeneous settings.
Reward Shaping. Potential-Based Reward Shaping (PBRS) (Ng et al., 1999) shares with DG-PG the property of preserving optimal policies, but the two operate at different levels. PBRS reshapes the scalar reward via a potential function to provide denser temporal feedback, accelerating single-agent exploration. However, the shaped reward still aggregates all agents’ actions, so the resulting gradient estimator retains $\Theta(N)$ cross-agent variance. DG-PG instead operates at the gradient level, augmenting each agent’s policy gradient with a signal computed analytically from the reference model.
@@
Structural Priors in RL. Guided Policy Search (Levine and Koltun, 2013) and Residual RL (Johannink et al., 2019) incorporate domain knowledge by using pre-computed controllers as base policies or supervisory signals, but this couples the learned policy to the quality of the prior and can bias the solution. Physics-Informed Neural Networks (Raissi et al., 2019) embed governing equations directly into the loss function, but require exact PDEs and do not decompose across agents. DG-PG uses analytical models differently, constructing noise-free per-agent gradients of a guidance term that reduce gradient estimate variance without altering the optimization landscape. A systematic comparison is provided in Appendix A.
@@ 6.1 Setup
Environment. We evaluate on a heterogeneous cloud scheduling simulator with NN servers whose configurations are based on AWS instance types. The environment combines three sources of complexity (heterogeneous server efficiency, bimodal workloads with heavy tails, and non-stationary arrivals) that jointly violate the stationarity assumptions of the queueing-theoretic prior, testing robustness to model mismatch. The reward balances two competing objectives: latency (queue minimization) and energy efficiency (preferring high-$\eta_{\text{CPU}}$ servers), which cannot be simultaneously optimized by simple heuristics. Figure 1 illustrates these characteristics. We evaluate at scales $N\in\{5,10,20,50,100,200\}$. For each scale, we pre-generate 200 training scenarios and 40 held-out test scenarios with randomized server configurations and workload patterns. Full specifications are in Appendix H.
@@ 6.2 Hyperparameter Selection
The guidance weight $\alpha\in(0,1)$ controls the bias-variance trade-off in DG-PG. Larger $\alpha$ reduces gradient variance by relying more heavily on the analytical prior, while smaller $\alpha$ maintains closer alignment with the true reward signal. We conduct a systematic grid search over $\alpha\in\{0.2,0.4,0.6,0.8\}$ at $N=50$ using NN models, with validation at $N=20$.
@@ Results.
Figure 2 shows training curves at N=50N{=}50 for all tested $\alpha$ values. The results reveal a clear trade-off in which larger $\alpha$ accelerates early convergence, while smaller $\alpha$ achieves marginally superior final performance. By episode 200, $\alpha{=}0.2$ achieves the best mean training reward of 40.3-40.3, while $\alpha{=}0.8$ reaches 40.8-40.8, a gap of less than 1%. This trend is consistent at N=20N{=}20, where $\alpha{=}0.2$ achieves 39.9-39.9. The narrow spread across $\alpha$ values demonstrates that DG-PG is robust to the choice of $\alpha$, as the RL gradient compensates for model inaccuracy while benefiting from variance reduction at any tested setting.